OpenStudy (anonymous):

5x^2-9x-2=0 solve for x

OpenStudy (ghazi):

use \[D=\frac{ -b \pm \sqrt{(b)^2-4ac} }{ 2a }\] where a =5,b=-9 and c=-2

OpenStudy (anonymous):

I can't factor?

OpenStudy (ghazi):

no i guess you can't factor..you've to use this

OpenStudy (anonymous):

umm ok

OpenStudy (anonymous):

i hate using the quadratic formula

OpenStudy (ghazi):

but it's compulsion in this case

OpenStudy (anonymous):

yea ]

OpenStudy (anonymous):

can i use x^2-9x-2=0?

OpenStudy (ghazi):

no coefficient of x^2 is 5 ...you can't do that

hero (hero):

Why not just split the middle term to get: 5x^2 -10x + 1x - 2 = 0

OpenStudy (anonymous):

and do what?

hero (hero):

Factor the first two terms, then factor the last two terms to get: 5x(x-2) + 1(x-2) = 0

OpenStudy (anonymous):

so x=2, -1/2

OpenStudy (ghazi):

yes

hero (hero):

I saw that @ghazi

OpenStudy (anonymous):

no wait your wrong

OpenStudy (ghazi):

but i deleted it before somebody could have seen

OpenStudy (anonymous):

its -1/5 and 2

OpenStudy (ghazi):

yes that's correct